OpenStudy (anonymous):

substitution method 2x +5y =27 x=17 -3y

OpenStudy (holly00d1248):

2(17-3Y)+5Y=27

OpenStudy (holly00d1248):

then distribute

OpenStudy (anonymous):

distribute the y's

OpenStudy (holly00d1248):

no imean 2x17 and 2x-3Y

OpenStudy (anonymous):

2(34-6y)+5 =27

OpenStudy (holly00d1248):

onec you distribute it will look like this 34-6Y+5=27

OpenStudy (nothingwasthesame):

no babe the answer is 34 but thats it then it minus what ever else is in the parenthesses

OpenStudy (holly00d1248):

now get Y on one side

OpenStudy (holly00d1248):

so Y=2